A coil of inductive reactance 31Ω has a resistance of 8Ω. It is placed in series with a condenser of capacitive reactance 25Ω. The combination is connected to an a.c. source of 110V. The power factor of the circuit is ...

Grade:12

2 Answers

Arun
6 years ago
Power factor = R/Z
Where R is resistance and Z is impedance.
R = 8 ohm
Xc= 25 ohm
Xl = 31 ohm
Z = \sqrtR2 - (Xl --Xc)2
Z=\sqrt64 +36
Z = 10 ohm
Hence 
Power factor = 8/10 = 0.8
Kshirodra sahu
5 years ago
A choke coil has an impedance of 15 ohm and resistance 12 ohm find out the inductive seactor(xl) and inductance (l) of the coil if the frequency equal to 50hz also find the fower fectore